how do i write an equation that is parallel to y=-1/2x+4 that goes through these points (-3,5)

Found 2 solutions by TutorDelphia, mananth:
Answer by TutorDelphia(193)

parallel equations always have the same slope
since the equation you have is in slope intercept form we know the slope is -1/2
So far we know y=(-1/2)x+b
but we don't know b, the intercept so just plug in x and y from your point and solve for b
5=(-1/2)(-3)+b
5=3/2+b
10/2=3/2+b
7/2=b
y=(-1/2)x+7/2
As you can see our lines are parallel

Answer by mananth(16949)

y=-0.5x+4

Compare this equation with y=mx+b
slope m = -0.5
The slope of a line parallel to the above line will be the same
The slope of the required line will be -0.5
m= -1/2 ,point ( -3 , 5 )
Find b by plugging the values of m & the point in
y=mx+b
5 = 3/2 + b
b= 7/2
m= -1/2
Plug value of the slope and b in y = mx +b
The required equation isy = - 1/ 2x +7/2
m.ananth@hotmail.ca